### v3

# Optimal trajectories for Bayesian olfactory search in turbulent flows: the low information limit and beyond

Abstract
In turbulent flows, tracking the source of a passive scalar cue requires exploiting the limited information that can be gleaned from rare, stochastic encounters with the cue. When crafting a search policy, the most challenging and important decision is what to do in the absence of an encounter. In this work, we perform high-fidelity direct numerical simulations of a turbulent flow with a stationary source of tracer particles, and obtain quasi-optimal policies (in the sense of minimal average search time) with respect to the empirical encounter statistics. We study the trajectories under such policies and compare the results to those of the infotaxis heuristic. In the presence of a strong mean wind, the optimal motion in the absence of an encounter is zigzagging (akin to the well-known insect behavior “casting”) followed by a return to the starting location. The zigzag motion generates characteristic $t^{1/2}$ scaling of the rms displacement envelope. By passing to the limit where the probability of detection vanishes, we connect these results to the classical linear search problem and derive an estimate of the tail of the arrival time pdf as a stretched exponential, which agrees with Monte Carlo results. We also discuss what happens as the wind speed decreases.
Introduction. A wide variety of animals, from plankton [1] to fruit flies [2] to dogs [3], combine olfactory sensing of odors with complex search behaviors which enable them to track the sources of the odors [4]. A famous, striking example is the ability of moths to locate, starting from $\gtrsim 100$ m away, potential mates using a single sex pheromone cue to which they are sensitive via olfactory organs [5, 6, 7]. In many settings, including those relevant to flying insects [8], this olfactory search behavior occurs in a fully turbulent flow. Turbulent conditions are also relevant to chemical-sensing robots, which may be used to locate hazardous substances, fires or explosive devices [9, 10, 11, 12].
The consequences of turbulence in the olfactory search problem, by now well-studied [13, 14, 15, 4], cannot be overstated. Odors may be modeled as passive scalars; the turbulent transport of passive scalars exhibits a remarkable degree of spatiotemporal intermittency, and the concentration fluctuations exhibit strongly non-Gaussian statistics [16]. As odors are emitted by a stationary and continuous source and mix into the surrounding flow, they form a characteristic concentration profile called a plume, whose shape depends on details of the flow. Within the plume, the concentration fluctuates wildly on account of the turbulence, forming a complex and patchy landscape (see Fig. 1 for a visualization from simulation). If a searcher is sensitive to an odor above some detection threshold, its encounters with the odor are randomized and typically rare (except when the searcher is close to the source) [15]. Moreover, the odor typically manifests in isolated patches rather than a continuous field, rendering simple gradient-based search strategies effectively useless.
How should a searcher respond to these limited, random encounters in order to locate the source in minimal average time? Looking to biology can provide some inspiration. During upwind flight, a diverse array of flying insects exhibit a behavior called “casting,” which consists of broad zigzags perpendicular to the wind direction. Wind tunnel experiments suggest that this behavior is a very generic response to loss of contact with the odor cue [17, 18]. [In biological literature, “zigzagging” often refers to a specific insect behavior consisting of small-scale crosswind motion combined with upwind flight. We will frequently use the word “zigzagging” as simply a qualitative description of trajectories, without intending reference to this behavior.]
For robotics applications, this observation may be used to formulate biomimetic strategies, as in [14]. Alternatively, if the searcher is equipped with a model for the spatially-dependent likelihood of an encounter, the search problem may be formalized as a Bayesian partially-observable Markov decision process (POMDP) [19]. This formulation allows the optimal, i.e., minimum average time, search strategy to be described by a dynamic programming (Bellman) equation, which may be approximately solved [20, 21, 22].
The model likelihood is the key physics input into the problem. Previous work on the POMDP approach imposed a model likelihood ad hoc, with the effects of turbulence captured by an effective diffusivity only. The realism of such a model is highly dubious in light of the complexity of turbulent transport; it is certain to fail, for example, in the large wind regime where ballistic transport dominates. (In the absence of a good model, the search problem could instead be attacked by model-free reinforcement learning techniques, as in Refs. [23, 24].)
In order to maximize the realism of the model, in this work we study optimal trajectories with respect to likelihoods taken directly from simulation data. In particular, we examine the trajectories which result from the absence of an encounter; because encounters are rare, in a typical search, most of the search time will be spent in this low-information phase.
The optimal trajectories which we compute are complex and depend strongly on the strength of the mean wind. We focus on the case where the mean wind is strong: in this case the optimal motion is to zigzag upwind similarly to casting. This motion exhibits characteristic diffusive scaling in both directions. At some critical time, if the searcher has still not found the source, it should return to its starting location. We explain these results using a simple model that assumes encounters are very rare except close to the source. The diffusive scaling is explained by a reduction to the classical linear search problem and is used to predict the arrival time pdf.
Methods. From the POMDP point of view, the searcher is a Bayesian agent living on a gridworld which maintains a posterior probability density (or “belief”) $b(\bm{x})$ over the possible relative positions of the source with respect to the agent; critically, this relative position is unknown. The searcher makes an observation at every timestep, which we define as measuring if the local concentration is above or below some threshold $c_{\rm thr}$. [In accordance with usual POMDP practice, the observation rate $\nu_{\Omega}$ is the same as the action rate, i.e. $\nu_{\Omega}=v/\Delta x,$ where $v$ is the speed of the searcher and $\Delta x$ is the grid spacing. In all figures, we will express time in units of $\nu_{\Omega}^{-1}$ and distance in units of $\Delta x.$] The belief is updated in two ways. When the searcher moves (“takes an action” in POMDP jargon), the belief is transported according to $b(\bm{x})\to b(\bm{x}-\delta\bm{x}),$ where $\delta\bm{x}$ is the searcher’s displacement under the action. Secondly, when the searcher makes an observation $\Omega=\theta(c(\bm{x})-c_{\rm thr}),$ it updates its belief according to Bayes’ theorem [25]:
$b(\bm{x})\to b(\bm{x})p_{\Omega}(\bm{x})/Z,$ (1)
where $Z$ is the appropriate normalization constant and we have introduced the notation $p_{\Omega}(\bm{x})\equiv{\rm Pr}(\Omega|\bm{x}).$ The likelihood $p_{\Omega}(\bm{x})$ must therefore be assumed to be known to the searcher. In the case where the mean wind $U\hat{x}$ is strong compared to flow velocity fluctuations, it can be modelled as proposed in [15]. The Lagrangian particle motion is predominantly ballistic on the short times relevant to this limit, which facilitates analysis of the tail of the concentration distribution. In particular, the probability that the instantaneous concentration is higher than $c_{\rm thr}$ is
$p_{1}(\bm{x})\propto\theta(x)x^{-\alpha}\exp\left[-\left(\frac{Uy}{u_{\rm rms}x}\right)^{2}\right]\exp(-x/x_{D}),$ (2)
where $u_{\rm rms}$ is the typical fluctuation speed of the flow, $\theta$ is a Heaviside function, we have supposed the source is in $x=0$, $y$ is the distance from the plume centerline, $\alpha\geq 0$ is a free parameter that depends on the environment, and $x_{D}$ is a characteristic decay length which depends on $c_{\rm thr}$. This result is only valid sufficiently far from the source.
On the other hand, when the mean wind is small, long diffusive timescales become relevant to the Lagrangian particle motion and one instead expects
$p_{1}(\bm{x})\propto|\bm{x}|^{-1}\exp\left(-|\bm{x}|/\lambda\right)\exp(Ux/2D),$ (3)
with $D$ the effective (Taylor) diffusivity, $\lambda=\sqrt{D\tau/(1+U^{2}\tau/4D)},$ and $\tau=\tau(c_{\rm thr})$ is a typical time for a fluid puff, initially at the source, to grow large enough that the concentration is less than $c_{\rm thr}.$
Rather than impose one of these models directly, we used empirical models taken from simulations. We performed high-fidelity direct numerical simulations (DNS) of a turbulent flow (${\rm Re}_{\lambda}\simeq 150$), while tracking the positions of tracer particles which are emitted by a stationary point source. The system is advected by a uniform mean flow $U\hat{x};$ we repeated the simulation for $U/u_{\rm rms}\in\{0,1.2,2.5,4.9,7.4\}$. The tracers are taken as proxies for a passive scalar cue; this is a good approximation at high Péclet number. After coarse-graining the particle data onto a 2–D grid of about $\simeq 3250$ cells, we imposed a threshold concentration $c_{\rm thr}=200$ particles/cell and computed the likelihood $p_{1}(\bm{x})$ that this threshold is instantaneously exceeded (see Fig. 1). For the largest values of $U,$ the data are a good fit to Eq. 2 with $\alpha=1,$ and for the smallest values of $U$ the data fit Eq. 3 (see Fig. 3 of the Supplemental Material  [^1] See Supplemental Material at [hyperlink], which includes Refs. [43, 44], for detailed information about methods for the DNS and POMDP solution, as well as additional calculations, plots, and movies.).
Figure 1: Upper panel (a): snapshot from the DNS showing particles (blue) and the 2–D slab (green) on which the data are coarse-grained. Lower panel (b): likelihood $p_{1}(\bm{x})$ of an encounter obtained from the DNS for $U/u_{\rm rms}\simeq 7.4.$
For each flow, we then used the SARSOP algorithm [27] to extract a quasi-optimal search policy from the underlying dynamic programming equation [28, 19] (Supplemental Material, Eq. 3 [26]). To study trajectories, we generated large ensembles of trajectories via Monte Carlo trials; for comparison, we did the same for the popular heuristic policy “infotaxis” [29], which seeks to minimize the entropy of the posterior.
An important detail is that all trajectories begin with an encounter at time zero, modeling the fact that the searcher has no incentive to search in the absence of an encounter. This induces a prior $b_{0}(\bm{x})\propto p_{1}(\bm{x}).$ To ensure consistency, we begin each trajectory by randomly selecting a source position from the prior. We also constrain the searcher to move in two dimensions, which both eases the computational burden of the POMDP and simplifies the analysis of the results. An example quasi-optimal trajectory is shown in Fig. 2.
Figure 2: Snapshots from a trajectory produced from a quasi-optimal policy. The source is shown as a green star and the searcher as a red square. Each encounter is shown as a blue circle. The searcher’s belief $b(\bm{x})$ on the position of the source is shown in grayscale. The searcher found the source a couple timesteps after $t=75$.
Results. Our results can be organized and understood through the following principle: because encounters are rare, the most important optimization the searcher must perform is to decide how it should move when it does not smell anything, and the single most likely trajectory is the one generated in the absence of any encounters after time zero. We will call this deterministic trajectory the “no-hit trajectory.” To support this claim, we evaluated the probability $P_{\rm nh}(t)$ that, at time $t,$ the searcher has not encountered the cue since time 0, given that it has not yet found the source. This is plotted for the largest wind speed in the upper panel of Fig. 4; the results for other wind speeds are very similar and shown in Figs. 4–5 in the Supplemental Material [26]. For both quasi-optimal and infotactic motion, $P_{\rm nh}(t)\gtrsim 0.5$ for all times $t$, so the no-hit trajectory plays an outsize role.
Figure 3: Quasi-optimal no-hit trajectory, for $U/u_{\rm rms}\simeq 7.4$ The searcher is shown as a red square, the searcher’s belief is shown in grayscale, and the starting point is a blue circle.
Figure 4:
We find that when the mean wind is strong, the quasi-optimal no-hit trajectory begins by moving slowly upwind while zigzagging in the crosswind direction (Fig. 3). This motion strongly resembles casting. However, at some critical time, the searcher ceases this behavior and returns downwind to the vicinity of its starting position. From then on, the motion is dominated by upwind and downwind motion. Downwind returns have also been occasionally observed in wind-tunnel experiments after a long time without an encounter [30, 31], but this behavior is far less well-known. In contrast, infotaxis produces almost pure zigzagging motion in the absence of an encounter, and essentially never executes a downwind return.
Other examples of no-hit trajectories are supplied in Secs. VI C–D of the Supplemental Material [26]. As the wind speed decreases, the downwind return occurs earlier and earlier, and zigzagging quickly becomes irrelevant to the optimal motion. Instead, at moderate wind speed, a quasi-optimal searcher exhibits repeated up- and downwind motion that begins to resemble a complex, asymmetric spiral. We suggest that the curves tend to approximately lie on isolines of the prior, which would imply the trajectories are essentially performing gradient ascent of the likelihood profile. When $U$ and $u_{\rm rms}$ are comparable, the problem has very little symmetry and, accordingly, the trajectories are especially complex. At $U=0,$ both the quasi-optimal and infotaxis trajectories resemble Archimedean spirals, consistent with previous studies [29, 32, 20].
These behaviors generate clear signatures in the mean squared displacement (MSD) of the searcher, $\langle\Delta x(t)^{2}\rangle\equiv\langle(x(t)-x(0))^{2}\rangle$ and $\langle\Delta y(t)^{2}\rangle\equiv\langle(y(t)-y(0))^{2}\rangle,$ of the searcher, which we plot in Fig. 4 (lower panels). These are averaged over an ensemble of $10^{4}$ trajectories, and compared with the squared displacements resulting from a no-hit trajectory. The results are very similar, suggesting that motion in absence of an encounter is dominating the MSD signal. For the larger values of $U,$ the MSD in $y$ exhibits fast oscillations, the amplitude of which grows like $t$. The MSD in $x$ also grows approximately linearly. These results imply that the rms displacement envelopes in both $x$ and $y$ exhibit diffusive scaling: $x,y\sim t^{1/2}$.
[…]
Low-information limit. The results at large wind speed can be understood by means of a simplified model. In particular, we again assume the searcher initializes its belief as the prior $b_{0}(\bm{x})$ induced by the encounter at time zero, i.e., proportional to Eq. 2 for large enough $x.$ However, consistent with the dominance of the no-hit trajectory, we assume successive encounters only occur when the searcher is close to the source—the likelihood is very close to unity in a strip directly downwind of the source (see yellow region in Fig. 1, lower panel), which is where $\langle c\rangle>c_{\rm thr}.$ Empirically, this strip has negligible length except when $U$ is large;
Explicitly, we model the likelihood of an encounter after time $t=0$ as
$p_{1}(\bm{x})=\begin{cases}1,&0<x<\ell^{*},\,y=0\\ 0,&\textrm{otherwise,}\end{cases}$ (4)
where $x$ and $y$ are, respectively the upwind and crosswind position of the source with respect to the searcher, and $\ell^{*}$ is the characteristic length of the strip. The searcher uses this likelihood to update its belief. We will call this simplified model the low-information limit. A similar model was used to describe sector search in Ref. [33].
In this limit, the result of the Bayesian update (Eq. 1) is that the searcher will simply zero out its prior in a $\ell^{*}$-neighborhood (and renormalize) as it moves (see Fig. 5, panel (a)). The optimal trajectory will then be the one that most efficiently explores the prior’s support, so the geometry of the prior plays an essential role.
Figure 5: Upper panel (a): in the simplified model, the prior decays as a function of $x$ and is only supported on a cone symmetric about the $x$-axis. The searcher (red circle, labeled “S”) sweeps out probability mass in the upwind direction as it moves, over a characteristic length $\ell^{*}.$ Lower left panel (b): a zigzag trajectory may be approximated by a sequence of crosswind motions at upwind distances $x_{k},$ and the decision problem reduces to selecting the $x_{k}.$ Lower right panel (c): when the searcher has swept out the cone up to an upwind distance $x_{R}\sim x_{D},$ the local value of the prior $b_{0}$ is comparable to that near the starting point and just outside the cone, $\bm{x}_{1}$. The searcher should return to $\bm{x}_{1}.$
We now determine the optimal trajectories in the low-information limit. Due to the Gaussian factor in Eq. 2, the prior will be strongly confined to a cone $C$ given by $|y|<ax$ with $a\sim u_{\rm rms}/U,$ so it is useful to approximate the confinement as exact: suppose the prior depends only on $x$ inside the cone, and it is zero outside (see Fig. 5, panel (a)). We seek to minimize the expected time of arrival
[…]
where $u\equiv y/x,$ $T(\cdot)$ is the time that the searcher first sweeps out a given point in space, and $p(x)\equiv xb_{0}(x).$
[…]
A zigzag trajectory can be (for large enough $x$) modeled by a sequence of purely crosswind moves at positions $x_{1},x_{2},\dots$ (see Fig. 5, panel (b)). A key observation is that the searcher should not move further upwind than $\ell^{*}$ during any crosswind traversal, i.e., we have the constraint $x_{k+1}-x_{k}\leq\ell^{*}$ for each $k$. Otherwise, there will be points not swept out by the trajectory, and because the trajectory is monotonic in $x,$ $J$ will be infinite.
In the absence of the inequality constraint, the decision problem of choosing the optimal sequence $\{x_{k}\}$ is exactly equivalent to the linear search problem (LSP), a well-studied decision problem which asks for the turning points of a trajectory on a line which minimizes the expected arrival time to a target drawn from some known probability density [35, 36, 37, 38]. This is shown in the Supplemental Material (Sec. V), after which we prove that if $p(x)$ decays exponentially or slower—as it does in our case—then $x_{k+1}-x_{k}$ asymptotically diverges for large $k$ [26]. Thus the inequality constraint must saturate far enough upwind, i.e., $x_{k+1}-x_{k}=\ell^{*},$ and one can show this gives rise to diffusive scaling as observed in the MSD data.
(ignoring numerical factors)
$x(t)\sim\sqrt{\frac{\ell^{*}vUt}{u_{\rm rms}}};\;y(t)\sim\sqrt{\frac{\ell^{*}vu_{\rm rms}t}{U}}\cos\sqrt{\frac{vUt}{\ell^{*}u_{\rm rms}}},$ (5)
with asymptotically constant speed $\sim v$ in an rms sense.
We therefore conclude that the zigzag motion with $t^{1/2}$ scaling is a consequence of the fact that (a) the likelihood decays exponentially with $x$ (and not faster), (b) the likelihood decays faster than exponential outside the cone, and (c) the likelihood is small outside of a strip in front of the source. However, the argument (in particular, point (b)) rests on the approximation $\exp[-(u_{\rm rms}y/Ux)^{2}]\ll\exp(-x/x_{D}).$ When the searcher reaches an upwind distance $x_{R}\sim x_{D},$ this approximation breaks down, and the local prior probability will be comparable to that just outside the cone, near the starting point (see Fig. 5, panel (c)). This means that there is no continued benefit to search upwind, and the searcher should instead return downwind—as we indeed observe in the quasi-optimal policy.
We cannot analytically predict the optimal trajectories at finite $U$ (and moreover the approximation Eq. 4 becomes untenable), but the low-information limit still grants some insights. Using the analysis of Ref. [15], one expects $x_{D}\sim u_{\rm rms}^{-1}(SU/c_{\rm thr})^{1/2}$ (where $S$ is the emission rate) when $U$ is large enough that ballistic single- and two-particle dispersion scaling still holds on relevant timescales. This is in basic agreement with our observation that downwind returns occur earlier as $U$ decreases. Eventually, as $U$ becomes comparable to $u_{\rm rms},$ diffusive motion begins to dominate trajectories of Lagrangian particles on timescales of interest. In this regime, the likelihood is no longer well-confined to a cone and gains some support even upwind of the source (this is visible in our simulations at $U/u_{\rm rms}\simeq 1.2$). This complexification of the prior geometry leads to concomitantly complex optimal trajectories (see Supplemental Material, Sec. VI A, C, D [26]).
Finally, a useful consequence of the $t^{1/2}$ scaling at large $U$ is that it allows us to predict the tail of the arrival time pdf analytically. In the low-information limit, the arrival time is dominated by the time of the first encounter after $t=0,$ which is the time it takes to reach a $\ell^{*}$-neighborhood of the source. Therefore, we can estimate the arrival time pdf $p(T)$ as
$p(T)=\int d^{2}\bm{x}\,b_{0}(\bm{x})\delta(T-T^{*}(\bm{x})),$ (6)
where $T^{*}(\bm{x})$ is the time that the optimal trajectory first reaches $\bm{x}.$ $b_{0}(\bm{x})$ is proportional to the likelihood, which is given by Eq. 2 for large $U,$ and we have from Eq. 5 that $T^{*}(x,y)\sim(u_{\rm rms}/U)(x^{2}/v\ell^{*}).$ This gives
$p(T)\propto T^{-\alpha/2}\exp\left(-k\sqrt{T}\right),$ (7)
with $k\sim x_{D}^{-1}\sqrt{Uv\ell^{*}/u_{\rm rms}},$ a new prediction.
In Fig. 6 we plot the pdf of $\sqrt{T}$ for $U/u_{\rm rms}\simeq 7.4$, where $T$ is the time of arrival. There is a broad range $0\leq T\lesssim 500$ where the pdf appears to decay exponentially, implying $p(T)\sim\exp(-k\sqrt{T})$ as predicted. The agreement with exponential decay is more precise for infotaxis; we could have anticipated this since infotaxis was observed to obey $t^{1/2}$ scaling for the entire trajectory, whereas this scaling breaks down for the quasi-optimal policy once it begins its downwind return.
Figure 6: Pdf of the square root of the arrival time $\sqrt{T}$, for $U/u_{\rm rms}\simeq 7.4,$ obtained using $10^{5}$ Monte Carlo trials.
In summary, our results suggest that the ubiquity of casting as a response to loss of plume contact may be due to its being the optimal no-hit trajectory in the limit of low information, in the presence of strong mean wind. The same limit makes nontrivial quantitative predictions about the temporal scaling of the trajectories and the arrival time pdfs, which agree with results obtained using quasi-optimal policies extracted from the dynamic programming equation.
Acknowledgements.
We thank Aurore Loisy, Massimo Cencini, Lorenzo Piro, and Fabio Bonaccorso for fruitful discussions and acknowledge useful interactions at the 2023 Les Houches summer school on turbulence. This work received funding from the European Union’s Horizon 2020 Program under grant agreement No. 882340. We also acknowledge financial support under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 published on 2.2.2022 by the Italian Ministry of University and Research (MUR), funded by the European Union – NextGenerationEU– Project Title Equations informed and data-driven approaches for collective optimal search in complex flows (CO-SEARCH), Contract 202249Z89M. – CUP B53D23003920006 and E53D23001610006.